OpenStudy (timaashorty):

write an equation in slope-intercept form for a line where slope = 0, and that passes through G(3,2)

OpenStudy (anonymous):

use this: y-y1=m(x-x1)

OpenStudy (anonymous):

you know that the slope which is the gradient m= 0

OpenStudy (anonymous):

and you've been given the x and y coordiinates

OpenStudy (anonymous):

all you have to do now is substitute the values for m, x1 and y1

OpenStudy (timaashorty):

Can you help me do that?

OpenStudy (timaashorty):

@crystal1_1

OpenStudy (anonymous):

ok so m=0 x1=3 y1=2

OpenStudy (anonymous):

the formula for the equation is y-y1=m(x-x1) therefore, y-2=0(x-3) y-2= 0timesx-0times3 y-2=0 y=2

OpenStudy (anonymous):

i think thats how you do it. I've never heard of slope intercept form but i'm assuming it's the same as gradient intercept.

OpenStudy (timaashorty):

Therefore it's y-2 = 0 (x-3) ?

OpenStudy (timaashorty):

Wait now I see what you did. I think you did it just right. thank you!